### v2

# Lifelong Learning of Video Diffusion Models From a Single Video Stream

Abstract
This work demonstrates that training autoregressive video diffusion models from a single, continuous video stream is not only possible but remarkably can also be competitive with standard offline training approaches given the same number of gradient steps. Our demonstration further reveals that this main result can be achieved using experience replay that only retains a subset of the preceding video stream. We also contribute three new single video generative modeling datasets suitable for evaluating lifelong video model learning: Lifelong Bouncing Balls, Lifelong 3D Maze, and Lifelong PLAICraft. Each dataset contains over a million consecutive frames from a synthetic environment of increasing complexity. [^1] Code and datasets will be released on acceptance.
1 Introduction
There are a plethora of names – lifelong learning, continual learning, streaming inference – given to a central desideratum of artificial intelligence (AI) systems: the ability to learn from a single continuous autocorrelated stream of data. However named, our community has long sought models and algorithms that learn in a fundamentally human way; from birth to death, learning as we live.
Modern AI systems do not learn in this manner but instead rely on a highly effective compromise: stochastic gradient descent from data streams made up of independently and identically distributed (i.i.d.) samples. Language models [71, 53], world models [33, 34] , and video models [35, 40, 5] are trained on random batches of short temporally-correlated segments, an approach that preserves short-range autocorrelations while approximating i.i.d.-ness through permutations. While effective, this approximately-i.i.d. learning paradigm does not offer a satisfying mechanism for updating models when new data arrive. Existing model updating techniques that are often used in practice – training from a checkpoint on the union of old and new data [3], fine-tuning on just the new data [80], or training completely anew from scratch [60] – do not operate in the training regime from which humans learn and suffer from problems such as high computation cost and forgetting [75].
Alternatively, SGD on autocorrelated data streams is considered by some to be a viable candidate for human-like lifelong learning [49]. While there is a raft of work indicating that gradient-based learning on autocorrelated data is possible [28, 43, 57], folk wisdom maintains that this learning setup is hard, impractical, and prone to failure, especially for deep networks. Evidence of these beliefs can be found throughout the literature. The optimization community has developed numerous mechanisms for alleviating the effects of data stream temporal dependencies [46, 31, 19], suggesting the existence of problems with learning from autocorrelated data streams. Additionally, the Bayesian learning community has proposed numerous continual learning approaches through posterior updating [6, 13, 54, 7], though these theoretically sound approaches lack in practical scaled results. These bodies of work suggest a strong desire towards avoiding non-i.i.d. learning, making it reasonable to assume that lifelong learning of video models from a single autocorrelated video stream might be difficult if not impossible.
The primary contribution of this paper is simple yet profound. We empirically demonstrate that diffusion-based video models can be learned in a lifelong manner from a single autocorrelated video stream, obtaining comparable performance to standard (offline) i.i.d. training. To the best of our knowledge, this is the first demonstration of non-i.i.d. training of video diffusion models. Remarkably, successful lifelong learning does not require a complex setup. We find that a minimal set of established continual learning techniques, such as experience replay with limited memory, is sufficient to make lifelong- and i.i.d.-trained models indistinguishable in their performance given the same numbers of gradient steps and batch size.
(a) Lifelong Bouncing Balls frames subsampled from 5 second long videos.
(b) Lifelong 3D Maze frames subsampled from 2.5 second long videos.
(c) Lifelong PLAICraft frames subsampled from 4 second long videos.
Figure 1: Ground truth video frames (1st row of each subfigure), offline learned models’ generated videos (2nd and 3rd rows), and lifelong learned models’ generated video frames (4th and 5th rows) for our datasets. The left two columns highlighted in red show a subset of frames used to condition the model. Videos generated by lifelong learned models trained with experience replay are diverse, visually plausible, and crucially indistinguishable from those of offline learned models in quality.
As a secondary contribution, we introduce three novel lifelong learning video datasets with varying levels of temporal correlation, data repetitiveness, rare events, and non-stationarity. Even under an academic computational budget, we observe stable learning and reliable short-range extrapolations on all three benchmarks. As video modeling is a component of world modeling, our findings have the potential to open up new life-like approaches to learning, planning, and control in embodied AI agents.
2 Background
Video Diffusion Models
Video Diffusion Models (VDMs) are a class of generative models capable of synthesizing high-quality, temporally consistent videos [78, 40, 35, 41, 32, 10, 9, 14]. Rooted in the principles of denoising diffusion, video diffusion models progressively refine random noise into coherent video frames across a series of iterative denoising steps. Extending image diffusion models to videos requires capturing temporal dependencies between frames while preserving single-frame quality. These requirements present unique challenges, including the computational overhead of processing a large number of frames as well as developing architectures that can capture both spatial and temporal coherence efficiently. To handle long-duration videos, models learn the conditional probability of new frames given previously generated frames and then generate videos in an auto-regressive manner [35, 24].
Lifelong Learning
The goal of lifelong or continual learning is enabling models to continuously learn from new data with minimal forgetting of what was learned before [23, 79, 73, 81]. Approaches for lifelong learning include using regularization terms to penalize changes to parts of the network that encode previously learned information [45, 85, 48], improving the plasticity or the stability of the optimization algorithm [25, 38, 82], and enforcing the encoding of different tasks in mostly non-overlapping or orthogonal parts of the model [64, 66, 84]. Another popular approach is replay, whereby samples representative of past data are revisited by the model along with the current training data [61]. Such replayed samples can be obtained from a generative model [67], but often they are sampled from a memory buffer containing past training data, an approach referred to as experience replay [20, 62, 15, 2].
Figure 2: Video diffusion model lifelong learning on a single video stream with $K\!=\!4$. At training step $t$, the model conditions on the frames in the first half of its context window (red) and learns to denoise the frames in the second half of its context window (blue). At training step $t+1$, the model’s context window shifts by one video frame, and the same procedure repeats indefinitely.
Most work on neural network lifelong learning has focused on the highly simplified problem setting where a classification model is trained on a sequence of non-overlapping tasks, with each task seen only once. Often the model can train on each new task until convergence (referred to as the offline setting), although some works only allow a single pass over the data of each task (the online setting) [1, 22]. Beyond classification, lifelong learning research has explored the continual training of generative models, including generative adversarial networks, variational auto-encoders and diffusion models for image generation [86, 47, 29, 69]. However, as far as we are aware, no work thus far has explored lifelong learning of video diffusion models. The lifelong learning community has also shown interest in moving beyond learning in a strictly task-based manner. To create data streams with more complex temporal correlations, blurry task boundaries [4, 52] and repetition of previously seen concepts [37] have been used. Benchmarks for lifelong learning have also been constructed based on data from autonomous driving [74], using image datasets collected through time [11] and by concatenating thousands of short videos [18].
3 Datasets
Exploring the possibility of learning video models in a lifelong fashion requires long continuous video streams. As discussed in Section 2, existing video datasets usually consist of many short videos that may or may not be temporally related. To that end, we introduce three new video datasets, each of which is constructed from a single video stream without semantic discontinuities. The three datasets – Lifelong Bouncing Balls, Lifelong 3D Maze, and Lifelong PLAICraft video datasets – vary in terms of complexity, stochasticity, temporal correlation, and degree of non-stationarity.
3.1 Lifelong Bouncing Balls
We introduce two versions of the Lifelong Bouncing Balls dataset: Version O for “original” and C for “changing”.
[…]
Lifelong Bouncing Balls (C) introduces non-stationarity on top of Lifelong Bouncing Balls (O). While the balls’ motion matches that of (O), the blue channel values of all ball colors increase over time at a constant rate (red, yellow, and green ball colors respectively become fuchsia, white, and aqua by the end of the video). The evaluation set contains frames with balls of all previously observed colors. Solving Lifelong Bouncing Balls (C) requires learning the deterministic ball dynamics and color transitions from a correlated video stream where some details are unrepeated.
3.2 Lifelong 3D Maze
The Lifelong 3D Maze dataset contains 1 million 64x64 RGB video frames for training ($\sim$14 hours long at 20FPS) and 100,000 video frames for evaluation. The video is a first-person view of an agent that navigates a randomly generated 3D maze. Whenever the agent solves a maze, the walls of the solved maze come down and the walls of an unseen maze rise, at which point the agent attempts to solve the new maze. The mazes contain various sparsely appearing objects, including polyhedral gray rocks that flip the agent upside down upon being touched. Since the maze states are partially observable and their transition dynamics are stochastic, the future frames cannot be perfectly predicted given the past frames. Solving Lifelong 3D Maze requires learning the first-person sensory inputs associated with navigating a stochastically generated 3D environment and modeling infrequent events from a largely repetitive and correlated video stream.
3.3 Lifelong PLAICraft
The Lifelong PLAICraft dataset contains 1.85 million 1280x768 RGB video frames for training ($\sim$54 hours long at 10FPS) and 600,000 video frames for evaluation. To make training computationally feasible, we encode these video frames using the Stable Diffusion perceptual encoder [63] to a shape of 4x160x96. The training video is a first-person view of an anonymous player with an in-game ID of “Alex” engaged in the PLAICraft project’s multiplayer Minecraft survival world [36]. The evaluation video is a first-person view of another anonymous player with an in-game ID of “Kyrie” who explores the same Minecraft world. The videos capture multiple continuous play sessions spanning several months within this shared multiplayer survival world, showcasing various biomes in all three world dimensions, mining, crafting activities, construction, mob fighting, and player-to-player interactions. The Minecraft world contains aspects that repeat (ex. day-night cycle, players visiting their homes) and do not repeat (ex. felled trees, player chat logs). Solving Lifelong PLAICraft requires learning a highly nonstationary environment that changes in multiple timescales from a single correlated video stream.
4 Model and Training Regime
Model
We use a U-Net video diffusion model with spatiotemporal attention introduced by Harvey et al. [35]. The video diffusion model’s U-Net operates on $K$ video frames at a time, where $K=10$ for Bouncing Balls and $K=20$ for the other two datasets. The first $K/2$ video frames are model inputs and the next $K/2$ video frames are prediction targets. During training, given the first $K/2$ uncorrupted video frames and the second $K/2$ video frames corrupted with Gaussian noise, the model regresses the values of the Gaussian noise [56]. During sampling, given the first $K/2$ clean video frames and the second $K/2$ video frames filled with Gaussian noise, the model iteratively “denoises” the second $K/2$ video frames until they resemble a plausible continuation of the first $K/2$ video frames.
Baseline: Offline (i.i.d.) Learning
As a baseline representative of standard video models, we train models in an Offline Learning regime. At each training step, the model has access to $K$ contiguous video frames sampled from a random point in the data stream. The resulting training data resembles i.i.d. segments of $K$ video frames, even though all frames originate from a single autocorrelated stream. When using batch size $N_{\text{batch}}>1$ (as is common), this i.i.d. sampling is repeated $N_{\text{batch}}$ times.
Lifelong Learning
While the Offline Learning regime randomly permutes the data during training, the Lifelong Learning regime presents data to the model in order. This setup reflects the desiderata outlined in Section 1, where the model learns from a real-time data stream. At training step $t$, the model receives the $t$-th to $t\!+\!K$-th video frames of the (single) video that defines the dataset. After the model has taken a gradient step, the model receives the next set of frames ($t+1$ to $t\!+\!K\!+\!1$).
Train Stream Test Stream
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 5.5 ±0.2 6.2e-5 ±2e-6 1.63 ±0.03 0.005 ±0.001 5.4 ±0.1 6.2e-5 ±2e-6 1.58 ±0.04 0.005 ±0.0
Lifelong Learning 5.0 ±0.1 6.3e-5 ±1e-6 1.71 ±0.05 0.004 ±0.0 5.0 ±0.2 6.2e-5 ±1e-6 1.68 ±0.01 0.004 ±0.0
Table 1: Lifelong Bouncing Balls (O) performance metrics (lower is better for all) computed across two training and three sampling seeds.
Train Stream Test Stream
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 6.9 ±0.3 6.5e-5 ±2e-6 1.92 ±0.09 0.007 ±0.002 8.0 ±0.5 6.5e-5 ±2e-6 2.01 ±0.09 0.006 ±0.001
Lifelong Learning 6.1 ±0.3 7.4e-5 ±0.0 1.89 ±0.03 0.005 ±0.0 6.7 ±0.3 7.5e-5 ±1e-6 1.96 ±0.01 0.005 ±0.001
Table 2: Lifelong Bouncing Balls (C) performance metrics (lower is better for all) computed across two training and three sampling seeds.
This setup is compatible with numerous techniques proposed by the continual learning community. To demonstrate this compatibility, we augment the online stream with a minimal experience replay buffer [20] that retains $N_{\text{buffer}}$ past video subsequences chosen through reservoir sampling [77]. $N_{\text{buffer}}$ is typically chosen to be a small fraction of the total data stream size ($\approx 5-20\%$). At training step $t$, the model trains on the $t$-th to $t\!+\!K$-th frames from the data stream as well as $N_{\text{batch}}\!-\!1$ random replay buffer sequences, resulting in the same batch size as standard Offline Learning. We emphasize that replay is a relatively simple and well-established technique that could be substituted with any combination of other continual learning methods. When no replay buffer is used, we fill each training batch with $N_{\text{batch}}$ replications of the $t$-th to $t\!+\!K$-th frames. The diffusion model training procedure samples different values of Gaussian noise for each of these replications, resulting in a reduced-variance estimate of the video diffusion model loss [40] (this can also be done for experience replay to prioritize learning new data).
5 Experiments
This section qualitatively and quantitatively investigates the capabilities of Offline Trained and Lifelong Learned video diffusion models. The learning algorithms use the same batch size and number of gradient steps, equalizing the amount of computation and runtime memory. We refer the readers to Appendix B for additional experiment details.
5.1 Lifelong Bouncing Balls.
As discussed in Section 3, the Bouncing Balls dataset features repetitive and deterministic dynamics, with stationary (Lifelong Bouncing Balls (O)) and non-stationary (Lifelong Bouncing Balls (C)) color dynamics.
Setup
The models for Lifelong Bouncing Balls (O) and Lifelong Bouncing Balls (C) respectively have 8 and 74 million parameters. The models are trained with a batch size of 2 and are evaluated using 45 frames that they generate conditioned on 5 ground truth frames. Lifelong Learning retains 5 percent of the video stream’s frames in the replay buffer. We measure the sampled videos’ perceptual and temporal coherence using FVD [72] and report the model loss to measure learning progress. We also measure the models’ understanding of the ball movement using minADE [59] and of the color transition using our ColorKL metric (Appendix A), extracting the ball positions and colors from the frames using 2D convolutions.
Train Stream Test Stream
Method FVD KVD Loss FVD KVD Loss
Offline Learning 93.5 ±1.3 66.6 ±0.9 0.0056 ±0.0 89.2 ±1.0 62.7 ±0.8 0.0056 ±0.0
Lifelong Learning 98.0 ±3.2 62.8 ±2.7 0.0068 ±0.0001 98.8 ±2.4 62.3 ±1.5 0.0059 ±0.0
Table 3: Lifelong 3D Maze performance metrics (lower is better for all) computed across two training and three sampling seeds.
Result
Figure 3: Sample comparison for Lifelong Bouncing Balls (C) where ball colors have changed from red/yellow/green. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(a) and Fig. 3. For Lifelong Bouncing Balls (O), Offline Learning and Lifelong Learning produce models that generate visually compelling ball colors and trajectories, correctly handling different ball-to-ball and ball-to-wall collisions. While no model perfectly recovers the deterministic ground truth trajectories, the videos generated by offline and lifelong learned models are not perceptibly qualitatively different. Because the model’s context size $K=10$ is too small to always condition on two past ball bounces, the models do not perfectly capture the color transition from red to yellow and green. However, the color transition from yellow to red and green to red are captured well, and the models probabilistically transition from red to yellow or green – the best possible modeling choice given the limited context size. Perhaps surprisingly, we observe similar behaviors for Lifelong Bouncing Balls (C). Despite the non-stationary color changes, both learning algorithms generate realistic ball trajectories and colors, regardless of whether we test on frames from the beginning or the end of the video stream.
The quantitative results in Tab. 1 and Tab. 2 mirror the qualitative results. Offline Learning and Lifelong Learning perform comparably on both the stationary Lifelong Bouncing Balls (O) and non-stationary Lifelong Bouncing Balls (C) datasets. Overall, we find that the presence of temporal correlation and the degree of non-repetitiveness do not pose a significant learning challenge to lifelong learning of video diffusion models with Lifelong Learning when the video stream is relatively simple.
Train Stream Test Stream
Method FVD KVD Loss FVD KVD Loss
Offline Learning 123.2 ±1.1 10.6 ±0.3 0.0330 191.8 ±1.6 14.9 ±0.3 0.0403
Lifelong Learning 107.3 ±1.2 8.3 ±0.2 0.0334 170.7 ±1.4 12.0 ±0.6 0.0408
Table 4: Lifelong PLAICraft performance metrics (lower is better for all) computed across one training and three sampling seeds.
5.2 Lifelong 3D Maze.
The Lifelong 3D Maze dataset – a first-person video stream from a partially observable navigation environment – is a step up in complexity from Lifelong Bouncing Balls. The frames in the video stream are renderings of a 3D space, whereas in Lifelong Bouncing Balls frames consist of 2D objects against a black background. Moreover, as discussed in Sec. 3, the 3D Maze data stream features randomly generated surroundings as well as rare events.
Setup
The diffusion models trained in this section have 78 million parameters. The models are trained with a batch size of 4 and are evaluated by generating 40 frames conditioned on 10 ground truth frames. Lifelong Learning retains 5 percent of the video stream’s frames in the replay buffer. We measure the sampled videos’ perceptual and temporal coherence using FVD and KVD [8] using the I3D network [17] from the FVD calculation. We also report the model loss to measure learning progress.
Result
Figure 4: Sample comparison for Lifelong 3D Maze’s rising walls rare event. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(b) and Fig. 4. Both Offline Learning and Lifelong Learning models generate visually compelling maze trajectories, successfully modeling sparsely occurring event sequences such as camera inversion and the rising of the maze walls. While the models occasionally make mistakes such as deforming polyhedral gray rock during camera inversion, these mistakes occur for both models.
Quantitative results appear in Tab. 3. Similar to the qualitative results, the performance metrics are comparable for Offline Learning and Lifelong Learning. Overall, lifelong learned models can capture information on rare events when learning from correlated video frames from a stochastically generated environment, despite heavy data imbalance between rare events and regular maze traversal frames.
5.3 Lifelong PLAICraft.
The Lifelong PLAICraft dataset is the most challenging of the three datasets. Like Lifelong 3D Maze, it is a continual video stream from a partially observable environment. However, the world is more complex with dynamics resulting from interactions between multiple agents.
Setup
The diffusion models trained in this section have 80 million parameters. Lifelong Learning retains 20 percent of the video stream’s frames in the replay buffer. The models are trained with a batch size of 8, of which two elements are the current timestep video frame subsequence, and are evaluated using 10 frames that they generate conditioned on 10 ground truth frames. We measure the sampled videos’ perceptual and temporal coherence using FVD and KVD, and report the model loss to measure learning progress.
Result
Figure 5: Sample comparison for Lifelong PLAICraft’s carrot harvesting sequence. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively.
Qualitative results appear in Fig. 1(c) and Fig. 5. We find that both Offline Learning and Lifelong Learning models capture perceptual details about Minecraft video frames despite having only 80 million parameters and 50 hours of gameplay data to work with. Objects present in every gameplay frame – player name, item bar, and the equipped item – are consistently included in all Offline Learning and Lifelong Learning generations. Compared to the other two datasets, both learning paradigms are less successful at capturing temporal correlations between frames. We hypothesize that a larger model size and careful hyperparameter tuning will bring the sample quality closer to that of the Lifelong Bouncing Balls and Lifelong 3D Mazes models. Nevertheless, the generated video frames from Offline Learning and Lifelong Learning models are not perceptually distinguishable (Appendix G contains additional results).
Quantitative results appear in Tab. 4. On both training and test streams, Offline Learning and Lifelong Learning perform comparably. To summarize, we find that Lifelong Learning can perform comparably to Offline Learning with minimal hyperparameter tuning, even on data streams as complex as Lifelong PLAICraft.
5.4 Discussion
Given the strong performance of Lifelong Learning, it is natural to wonder if performance could be improved (or worsened) by strengthening (or weakening) the continual learning techniques used as part of the learning paradigm. To that end, in Appendix D we repeat all Lifelong Learning experiments without the use of any replay buffer. Lifelong Learning still obtains competitive performance on repetitive datasets like Lifelong Bouncing Balls (O) and Lifelong 3D Maze. However, there is a more significant performance degradation on the non-repetitive Lifelong Bouncing Balls (C) dataset as well as Lifelong PLAICraft due to catastrophic forgetting.
At the same time, increasing the size of the replay buffer does not result in any significant performance improvement. In Appendix D we see that having access to an unlimited replay buffer – capable of storing all past frames – offers no significant performance improvement over the minimal replay buffer used in this section for all experiments. Taken together, these results suggest that offline performance can be achieved with very little continual learning methodology. We hypothesize that our minimal replay buffer may even be overkill, and simpler regularizations may also yield competitive Lifelong Learned models.
Lastly, to further emphasize that the strong performance of Lifelong Learning is not a byproduct of the generative modeling tasks being unchallenging, we plot the change in the video diffusion models’ test stream quantitative metrics during training for the easiest and hardest datasets, Lifelong Bouncing Balls (O) and Lifelong PLAICraft, in Fig. 6 (similar figures for all datasets are presented in Appendix E). Both Offline and Lifelong Learned models continue to improve performance as training progresses, indicating that the generative modeling tasks are not easy enough to be mastered after the models train on a moderate number of video frames.
(a) Lifelong Bouncing Balls (O) dataset train-time test stream performance metrics.
(b) Lifelong PLAICraft dataset train-time test stream performance metrics.
Figure 6: The evolution of test stream performance metrics during training at different training iterations. The plots show the improvement in model quality as online training progresses. Error bars reflect standard error computed across one training and three sampling seeds.
6 Related Work
Our work is closely related to continual generative modeling [55, 58, 51, 69, 83, 16, 21], in particular to methods that focus on diffusion models and video generative models. Prior work on lifelong learning of diffusion models [51, 69, 83] focuses on image modeling under the task-based continual learning setup where data grouped into disjoint tasks arrives in large batches. In contrast, our work focuses on lifelong learning of video diffusion models capable of capturing temporal correlations in video frames by learning from a single video stream. Prior work on lifelong learning of video generative models [16, 21] continually learns VAE-based models that can generate future frames again in the task-based continual learning setup. Unlike our setup, these methods assume the presence of rigid task boundaries and the ability to train to convergence on large data batches. In contrast, our lifelong learning datasets do not have notions of tasks, and our models only have access to data as they appear in the stream.
Our work is also related to online learning of sequence processing models [87, 18, 12, 50]. Notably, Carreira et al. [18] learns predictive video models from the concatenation of loosely related videos. In contrast, our work learns generative video models. Liu et al. [50] learns a linear regression-based world model in an online fashion, an approach that we note is not amenable to high-dimensional videos.
Lastly, several recent works have, like us, introduced video datasets for continual learning. Villa et al. [76], Tang et al. [70]’s datasets contain many short videos that can be learned under a task-based continual learning setup. Carreira et al. [18]’s datasets construct one very long data stream by concatenating multiple short-to-medium length videos, but their data are not publicly available. Singh et al. [68]’s dataset, while not originally developed for continual learning, contains a series of short-to-medium-length videos of the real world that were collected via Google Glass. We note that their dataset has many semantic discontinuities, whereas our datasets come from a continuous stream under a single environment.
7 Conclusion
This paper explores lifelong learning of diffusion models from video streams, which has never been investigated to the best of our knowledge. We establish the feasibility of learning video diffusion models from a single autocorrelated video stream in a lifelong fashion. Our most successful learning paradigm is remarkably simple and uses no specialty techniques beyond a minimal replay buffer. To promote further research into this area, we introduce three datasets that test how different video stream characteristics affect model and lifelong learning algorithm performance.
The ability to effectively train video diffusion models from a continuous data stream holds significant promise, particularly for world model learning [27, 42, 30] and compute-efficient foundation model adaptation [69]. Our findings suggest that learning generative models from strongly correlated data streams may not be as difficult as previously assumed, at least for certain data streams and model classes, echoing results from Bornschein et al. [12] and Carreira et al. [18] on discriminative model lifelong learning. We expect sample quality to improve significantly with data and compute scale and also with more advanced model architectures, optimization techniques, and continual learning algorithms. Further analysis of lifelong learned video diffusion model behavior, compute-efficient video model adaptation methods, and lifelong learning video models on more complex video streams are promising avenues of future inquiry.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs Program, Inverted AI, MITACS, the Department of Energy through Lawrence Berkeley National Laboratory, and Google. This research was enabled in part by technical support and computational resources provided by the Digital Research Alliance of Canada Compute Canada (alliancecan.ca), the Advanced Research Computing at the University of British Columbia (arc.ubc.ca), and Amazon.
Supplementary Material
Appendix A The ColorKL Metric
The ColorKL metric measures how faithful a model-generated video’s color transition statistics are to the ground truth color transition statistics at the dataset level. It is defined as
$\frac{1}{|\mathcal{C}|}\sum_{c_{old}\in\mathcal{C}}D_{\mathrm{KL}}\left(p_{*}(c_{new}|c_{old})\;\|\ p_{model}(c_{new}|c_{old})\right)$
where $D_{\mathrm{KL}}$ is the KL divergence, $\mathcal{C}=\{red,yellow,green\}$ is the set of possible ball colors, $p_{*}(c_{new}|c_{old})$ is the ground truth probability for a ball to switch colors from $c_{old}$ to $c_{new}$, and $p_{model}$ is the empirical probability of color transitions from $c_{old}$ to $c_{new}$ in the model-generated video dataset. We note that if the model’s context window is long enough to always capture two past ball bounces for all balls, the ball colors can be deterministically predicted. If the model’s context is not long enough, as is the case in our experiments, the ground truth transition probabilities $p_{*}$ state that the balls always transition from yellow to red and green to red but have a 50/50 chance of transitioning from red to yellow or green.
Appendix B Additional Experiment Details
We employ the U-Net video diffusion architecture from Harvey et al. [35], which in turn is based on the U-Net image diffusion architecture of Nichol and Dhariwal [56]. We train the model using the epsilon prediction objective [39] and generate video samples during test time using Euler’s method sampler of Karras et al. [44].
The minimum average displacement error (minADE) metric is computed by selecting the trajectory with the lowest average displacement error from 3 sampled trajectories for each evaluation video subsequence. The ColorKL metric is computed by tallying the transition statistics from 3 sampled trajectories for each evaluation video subsequence and computing the KL divergence of this empirical distribution with the ground truth transition statistics. The loss metric is calculated by uniformly sampling 10 noise levels for all evaluation set samples per model checkpoint, computing the diffusion MSE loss, and averaging the results.
As computing all metrics on the entirety of the video streams is prohibitively expensive, we select 1,000 video subsequences from the evaluation stream and calculate the metrics on those video frames for all reported metrics. All training and sampling seeds compute the metrics on the same set of video subsequences. For datasets without significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (O), Lifelong 3D Maze), we select the first 1,000 video subsequences from the evaluation video stream. For datasets with significant changes in video frame details throughout the video stream (Lifelong Bouncing Balls (C), Lifelong PLAICraft), we evenly select 1,000 video subsequences across the entire evaluation video stream with equal spacing.
Appendix C Additional Dataset Curation Details
The train and test set of Lifelong Bouncing Balls is generated by randomly sampling the two balls’ initial positions and velocities and deterministically updating them to satisfy the conservation of momentum.
The Lifelong 3D Maze dataset was created by concatenating two 10-hour-long Lifelong 3D Maze YouTube videos [26, 65] at the point where a maze from the first video is solved and the maze from the second video begins, as the maze screensaver teleports the agent to a newly generated maze on the completion of the previous maze. This concatenation ensures that there is no sudden switch in the environment dynamics (a slight perceptual switch exists as shown in Fig. 7).
(a) Example frame from video 1.
(b) Example frame from video 1.
(c) Example frame from video 2.
(d) Example frame from video 2.
Figure 7: Example video frames from first and second 10-hour-long YouTube videos used to construct Lifelong 3D Maze. While the frames are nearly identical, there are very subtle perceptual differences arising from how the uploaders recorded the two videos.
The Lifelong PLAICraft dataset was constructed by concatenating multiple consecutive gameplay session recordings from the same player of the PLAICraft server [36]. All game settings, such as viewing distance, shader configurations, and recording parameters (e.g., resolution and aspect ratio), were consistent across sessions. The dataset includes recordings of two players, Alex and Kyrie, who played in the same survival world alongside hundreds of other players. In their initial gameplay sessions, their starting locations were randomly assigned but could later be altered through travel or teleportation. As in a typical Minecraft survival world, the environment had no boundaries; landscapes were procedurally generated and remained unchanged once created. Player states, including inventories and spawn locations, were preserved between consecutive sessions.
Players’ locations and states were generally consistent across most consecutive gameplay sessions. In other words, if a player exited the game at a particular location and state in one session, they would typically resume from the same point in the next session. However, some aspects, such as the player’s viewing angle or environment, may be slightly different from reasons such as other players having built structures around the location the player logged off from or other technical reasons related to gameplay recording.
Kyrie’s first gameplay session began later than Alex’s, and there was some overlap in the locations they visited. Alex generally explored more areas and contributed extensively to building structures, whereas Kyrie primarily played in a village that Alex and other players had previously developed.
Appendix D Additional Quantitative Results
Train Stream Test Stream
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 5.5 ±0.2 6.2e-5 ±2e-6 1.63 ±0.03 0.005 ±0.001 5.4 ±0.1 6.2e-5 ±2e-6 1.58 ±0.04 0.005 ±0.0
Experience Replay 5.0 ±0.1 6.3e-5 ±1e-6 1.71 ±0.05 0.004 ±0.0 5.0 ±0.2 6.2e-5 ±1e-6 1.68 ±0.01 0.004 ±0.0
No Replay 5.3 ±0.2 6.7e-5 ±1e-6 1.79 ±0.01 0.005 ±0.0 5.4 ±0.1 6.6e-5 ±1e-6 1.75 ±0.02 0.004 ±0.001
Full Replay 5.3 ±0.2 6.2e-5 ±0.0 1.68 ±0.01 0.004 ±0.0 5.1 ±0.2 6.2e-5 ±0.0 1.62 ±0.01 0.004 ±0.001
Table 5: Lifelong Bouncing Balls (O) performance metrics. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds. Lower is better for all metrics.
Train Stream Test Stream
Method FVD Loss minADE ColorKL FVD Loss minADE ColorKL
Offline Learning 6.9 ±0.3 6.5e-5 ±2e-6 1.92 ±0.09 0.007 ±0.002 8.0 ±0.5 6.5e-5 ±2e-6 2.01 ±0.09 0.006 ±0.001
Experience Replay 6.1 ±0.3 7.4e-5 ±0.0 1.89 ±0.03 0.005 ±0.0 6.7 ±0.3 7.5e-5 ±1e-6 1.96 ±0.01 0.005 ±0.001
No Replay 343.5 ±1.6 2.3e-2 ±1e-3 2.59 ±0.09 0.022 ±0.0 341.1 ±1.8 2.2e-2 ±1e-3 2.68 ±0.1 0.023 ±0.001
Full Replay 5.6 ±0.3 7.9e-5 ±2e-6 1.99 ±0.01 0.005 ±0.001 6.3 ±0.4 7.9e-5 ±2e-6 2.03 ±0.0 0.006 ±0.002
Table 6: Lifelong Bouncing Balls (C) performance metrics. The left and right columns respectively denote training and test video stream results computed across two training and three sampling random seeds. Lower is better for all metrics.
This section presents the train and test video stream performances of offline training (Offline Learning), lifelong learning using experience replay with limited replay buffer size (Experience Replay) [^2] This configuration is referred to as Lifelong Learning in the main text’s Sec. 5., lifelong learning using experience replay with unlimited replay buffer size (Full Replay), and lifelong learning without the use of a replay buffer (No Replay) on all datasets. Additional qualitative results appear in Appendix G.
D.1 Lifelong Bouncing Balls
For Lifelong Bouncing Balls (O), all learning algorithms attain comparable results. However, for Lifelong Bouncing Balls (C), No Replay performs significantly worse on all metrics than the rest as No Replay’s generated ball colors are different from those in the conditioning frames when the conditioning frames contain ball colors observed early in the video stream (refer to Fig. 12 and Fig. 16). This suggests that when lifelong learning on video streams that contain non-repeating details, mechanisms that preserve past knowledge are necessary regardless of how simple the video streams might be. This conclusion is further supported by Fig. 9(b), where the fully trained No Replay model performs progressively worse on data from earlier in the training stream.
D.2 Lifelong 3D Maze
We observe that all learning algorithms attain comparable results on the test stream. However, No Replay performs significantly worse on the train stream metrics. This is in part due to the way the Lifelong 3D Maze dataset is constructed. The dataset is built by concatenating two 10-hour-long YouTube videos, and the test set only comprises unobserved frames from the second video. While the videos are nearly identical, there are very subtle perceptual differences (refer to Fig. 7). Because the training stream contains frames from the first video and the test stream contains frames from the second video, No Replay performs better on the test stream videos than on the train stream videos as it overfits to the perceptual details of the second video.
D.3 Lifelong PLAICraft
Train Stream Test Stream
Method FVD KVD Loss FVD KVD Loss
Offline Learning 93.5 ±1.3 66.6 ±0.9 0.0056 ±0.0 89.2 ±1.0 62.7 ±0.8 0.0056 ±0.0
Experience Replay 98.0 ±3.2 62.8 ±2.7 0.0068 ±0.0001 98.8 ±2.4 62.3 ±1.5 0.0059 ±0.0
No Replay 220.0 ±0.9 168.2 ±0.8 0.0086 ±0.0 90.3 ±1.0 54.6 ±0.6 0.0058 ±0.0
Full Replay 97.4 ±1.3 67.5 ±1.2 0.0055 ±0.0 92.3 ±2.1 63.2 ±1.6 0.0057 ±0.0
Table 7: Lifelong 3D Maze performance metrics. Results are computed across two training and three sampling random seeds. Lower is better for all metrics.
Train Stream Test Stream
Method FVD KVD Loss FVD KVD Loss
Offline Learning 123.2 ±1.1 10.6 ±0.3 0.0330 191.8 ±1.6 14.9 ±0.3 0.0403
Experience Replay 107.3 ±1.2 8.3 ±0.2 0.0334 170.7 ±1.4 12.0 ±0.6 0.0408
No Replay 400.7 ±3.1 27.3 ±0.3 0.0487 485.2 ±3.7 42.8 ±0.7 0.0530
Full Replay 121.8 ±1.8 10.0 ±0.3 0.0329 185.0 ±0.5 13.0 ±0.1 0.0398
Table 8: Lifelong PLAICraft performance metrics. Results are computed across one training and three sampling seeds. Lower is better for all metrics.
We observe that Offline Learning and experience replay-based methods perform similarly but No Replay performs significantly worse. Fig. 9(d) shows the final models’ performances on video frames at different training stream indices. To further analyze model behaviors, we plot the fully trained models’ performance metrics on video frames from different parts of the training stream. Interestingly, the relative performance difference between different training stream frame indices for the lifelong learned models, even No Replay, mirrors that of the offline-learned model. This suggests that video stream underfitting, not forgetting, is the primary challenge in effective learning of Lifelong PLAICraft as we can assume that the offline-learned model does not suffer from forgetting.
Appendix E Training Time Performance Metrics
(a) Lifelong Bouncing Balls (O) dataset.
(b) Lifelong Bouncing Balls (C) dataset.
(c) Lifelong 3D Maze dataset.
(d) Lifelong PLAICraft dataset.
Figure 8: Test stream performance metrics for model checkpoints at different training iterations. The plots show the improvement in model quality as online training progresses. All models generally improve the longer they are trained.
Appendix F Train Stream Performance Breakdown
(a) Lifelong Bouncing Balls (O) dataset.
(b) Lifelong Bouncing Balls (C) dataset.
(c) Lifelong 3D Maze dataset.
(d) Lifelong PLAICraft dataset.
Figure 9: Fully trained models’ performance metrics on video frames from different parts of the training stream. The plots show whether the final model performs better or worse on future frame prediction for frames earlier or later in the training stream. Each point of the plots is calculated using 1,000 consecutive video frame subsequences that succeed the corresponding train stream frame index.
Appendix G Additional Qualitative Results
G.1 Lifelong Bouncing Balls
G.1.1 Train Stream Examples
Figure 10: Lifelong Bouncing Balls (O) trajectories from a video early in the training stream.
Figure 11: Lifelong Bouncing Balls (O) trajectories from a video late in the training stream.
Figure 12: Lifelong Bouncing Balls (C) trajectories from a video early in the training stream.
Figure 13: Lifelong Bouncing Balls (C) trajectories from a video late in the training stream.
G.1.2 Test Stream Examples
Figure 14: Lifelong Bouncing Balls (O) trajectories from a video early in the test stream.
Figure 15: Lifelong Bouncing Balls (O) trajectories from a video late in the test stream.
Figure 16: Lifelong Bouncing Balls (C) trajectories from a video early in the test stream.
Figure 17: Lifelong Bouncing Balls (C) trajectories from a video late in the test stream.
G.2 Lifelong 3D Maze
G.2.1 Train Stream Examples
Figure 18: Lifelong 3D Maze crystal sequence from the training stream.
Figure 19: Lifelong 3D Maze painting scenario from the training stream.
G.2.2 Test Stream Examples
Figure 20: Lifelong 3D Maze crystal sequence from the test stream.
Figure 21: Lifelong 3D Maze painting scenario from the test stream.
G.3 Lifelong PLAICraft
G.3.1 Train Stream Examples
Figure 22: Crafting table scenario.
Figure 23: Chest scenario.
Figure 24: Looking at another player in front of the village scenario.
Figure 25: Walking forward on a plain scenario.
Figure 26: Walking forward by a mountain on the side scenario.
Figure 27: Sleeping scenario.
Figure 28: Walking forward in The End biome scenario.
G.3.2 Test Stream Examples
Figure 29: Crafting table scenario.
Figure 30: Chest scenario.
Figure 31: Walking forward on a plain area near some tall buildings scenario.
Figure 32: Swimming in an ocean with the f3 information menu on scenario.